if x represents the first of three consective even integers, express the sum if the three integers in the term x
Answer by checkley77(12569)

THE FIRST OF THREE CONSECUTIVE EVEN INTEGERS=X
THE NEXT EVEN INTEGER=(X+2)
THE THIRD EVEN INTEGER=(X+4)
THUS THE SUM IS:
SUM=X+(X+2)+(X+4).